The consumer's utility function for goods X and Y is U = 3X + 15Y. Good X is placed on the x-axis and good Y is placed on the y-axis. Which of the following statements is TRUE?

I. The marginal utility of good Y is 15.
II. The MRSXY = 5.
III. The consumer is always willing to trade away 5 units of good X for 1 unit of good Y.

A.

I, II, and III

B.

I and III

C.

I and II

D.

II only
Answer #1

As it is given that the consumer's utility function for goods X and Y is U = 3X + 15Y.

Good X is placed on the x-axis and good Y is placed on the y-axis.

MU=dU/dX

MUX=3(1)+0

=3

MUY=dU/dY

=0+15(1)

=15

MRS of XY=MUx/MUy

=3/15

=1/5

Since MU of X is 3 and MU of Y is 15. It means that the consumer is always willing to trade away 5 units of good X for 1 unit of good Y.

It means statement I and III are correct.
Hence option B is the correct answer.